# Probing soft signals of gravitational-wave memory with space-based interferometers

### 30 Mar 2026 v1

@@ VI Stochastic soft memory signals
A data stream of sufficiently long duration may contain multiple, or even a large number of, events with memory. In the latter case, the detector effectively observes a stochastic GW background (SGWB) of soft memory signals [99]. Even if the individual events cannot be measured, the SGWB may still be detectable if the event rate is sufficiently large. This could be the case, e.g., for the SGWB from compact binary scattering [36, 59]. The SGWB from compact binary mergers will also include a contribution from null memory, although it is expected to be sub-leading compared with the contribution from oscillatory GWs emitted by the low-frequency sources within the same population. In this section, we examine the signal PSD of an idealized SGWB of soft displacement-memory signals at the LISA-like detectors and assess its detectability. We will also briefly comment on the SGWB of soft velocity-memory signals.
@@
$h_{ij}(t,\mathbf{x})=\sum_{\lambda}\int_{-\infty}^{\infty}df\,d\Omega_{2}\,\tilde{h}_{\lambda}(f,\hat{\mathbf{k}})\,e_{ij}^{\lambda}(\hat{\mathbf{k}})\,e^{-2\pi if(t-\hat{\mathbf{k}}\cdot\mathbf{x})}.$ (66)
@@
$\left\langle\tilde{h}_{\lambda}(f,\hat{\mathbf{k}})\,\tilde{h}^{*}_{\lambda^{\prime}}(f^{\prime},\hat{\mathbf{k}}^{\prime})\right\rangle=\frac{S_{h}(f)}{16\pi}\,\delta(f-f^{\prime})\,\delta^{2}(\hat{\mathbf{k}},\hat{\mathbf{k}}^{\prime})\,\delta_{\lambda,\lambda^{\prime}},$ (67)
@@
For an ensemble of soft displacement-memory signals, the resulting $h_{ij}(t)$ is a superposition of step functions with different heights $\mathcal{H}$ and step positions $t_{*}$. Such a time-domain waveform is similar to a Brownian motion and is not wide-sense stationary. Nevertheless, the Welch PSD estimator turns out to be
@@
Now we consider an ensemble of mutually-independent soft displacement-memory signals with a uniform distribution of $\{\tau_{*},\lambda,\sin\beta,\Psi\}$ and a random distribution of $H$. In the present case, we refer to such a SGWB as “idealized”. A single event manifests in $h_{33}(t)$ as a step with height
$\mathcal{H}=\kappa H,\quad\kappa=-c^{2}_{\beta}\,c_{2\Psi}.$ (73)
It follows that $\langle\mathcal{H}\rangle=0$, and $\langle\mathcal{H}^{2}\rangle=\sigma_{\kappa}^{2}\left(\langle H\rangle^{2}+\sigma_{H}^{2}\right)$, with $\sigma_{\kappa}^{2}=4/15$. Eqs. (70) and (72) then yield
@@
where
$\Delta=\frac{2N}{T}\frac{\langle H\rangle^{2}+\sigma_{H}^{2}}{(2\pi f)^{2}}.$ (75)
The corresponding energy density $d\rho_{\text{gw}}/df$ [defined by Eq. (68)] is equivalent to a simple sum of the spectral energy flux density $\frac{1}{T}\frac{dE_{\text{gw}}}{R^{2}d\Omega_{2}df}=\frac{H^{2}}{8\pi T}$ associated with each memory jump. For example, if $H$ is distributed uniformly in $0,H_{\text{max}}]$, then $\langle H\rangle^{2}+\sigma_{H}^{2}=\frac{1}{3}H_{\text{max}}^{2}$. To verify this description, a time-domain simulation of the SGWB is performed and the first-generation TDI signal at LISA detector is computed using fastlisaresponse with the realistic ESA orbit. The SGWB is constructed from 4000 sources drawn from a uniform distribution of $\{H,\tau_{*},\lambda,\sin\beta,\Psi\}$. We find that the result agrees well with the analytical prediction to $S_{h}$ and the signal PSD given by Eq. (71), as shown in Fig. 22.
In principle, the signal PSD can be measured [45, 93], and the corresponding SNR squared is given by
@@ VII Summary and discussion
As predicted by general relativity, sources of high-frequency GWs can also produce soft signals in detectors sensitive only to low-frequency bands. Such signals take clean, universal forms, with the memory amplitudes as the only intrinsic parameters. We have investigated the prospects for detecting such signals with future space-based interferometers, through observations of individual events and the stochastic background. For a single event, the SNR in LISA, Taiji and TianQin can reach 10 for a displacement memory of amplitude $\sim 10^{-20}$. Such a signal could potentially be produced by compact binary scatterings, supernovae, Gamma-ray bursts, or certain dark transients. The SNR of a single event in BBO can reach 10 for a displacement memory of amplitude $\sim 10^{-24}$. Such signals are generically expected from stellar-mass compact binary mergers, due to the nonlinear nature of gravity. We examine the parameter-estimation precision for soft displacement-memory signals using simulated Bayesian analyses. To characterize the deviation of the GW spectrum from the zero-frequency limit or an exact soft waveform (e.g., in compact binary mergers), an effective correction factor can be introduced. We examine such corrections in realistic examples of compact binary scattering and quasi-circular BBH mergers, and find that the soft waveform can still serve as an effective template in these cases. A large number of unresolvable soft signals can also form a stochastic gravitational-wave background. For displacement memory, the resulting background is expected to follow a strain power spectral density $S_{h}\propto f^{-2}$. We evaluate the corresponding SNR of the signal power spectrum at the planned detectors considered in this paper. Our results indicate that soft signals from bursts with memory can be promising targets for space-based gravitational-wave detection.
@@ Appendix B Soft signals from dark radiation
Similar to gravitational radiation, memory can also appear in massless scalar and vector radiation [7]. If the scalar or vector field couples to the test masses in a GW interferometer, this memory can also lead to soft signals (or their mimickers, in the case of massive fields), [^6] Note that even in the absence of non-gravitational couplings, dark radiation generally produces a gravitational displacement memory, either null or ordinary depending on whether the field is massive. as will be shown below.
Consider a Klein–Gordon field $\phi$ and a Maxwell or Proca field $A_{a}$ sourced by point charges. The worldline action of a point particle with mass $m$, scalar charge $q_{\phi}$, and vector charge $q_{A}$ is given by
$S=\int(-m+q_{\phi}\phi)\,d\tau+\int q_{A}A_{a}dX^{a},$ (78)
@@
$\phi(u)$ $=\int_{-\infty}^{\infty}df\,\sum\tilde{\phi}_{l,m}(f)\,Y_{lm}\,e^{-2\pi ifu},$
$A^{0}(u)$ $=\int_{-\infty}^{\infty}df\sum\tilde{A}^{0}_{l,m}(f)\,Y_{lm}\,e^{-2\pi ifu},$
$A^{\text{L}}(u)$ $\equiv\mathbf{A}\cdot\hat{\mathbf{k}}=\int_{-\infty}^{\infty}df\sum\tilde{A}_{l,m}^{\text{L}}\,{}_{-1}Y_{lm}\,e^{-2\pi ifu},$
$A$ $\equiv\mathbf{A}\cdot(\mathbf{e}_{\Theta}-i\mathbf{e}_{\Phi})=\int_{-\infty}^{\infty}df\sum\tilde{A}_{l,m}\,{}_{-1}Y_{lm}\,e^{-2\pi ifu},$
(see Sec. III for the definition of $\{R,\Theta,\Phi\}$) where $u=t-v_{g}R$, with $v_{g}(f)=\sqrt{1-m^{2}/(2\pi f)^{2}}$ being the group velocity. The idealized frequency spectra of the non-vanishing modes for the 0PN scalar dipole radiation are
$\frac{\tilde{\phi}_{1,\pm 1}}{\frac{\nu\gamma aM}{R}}=\sqrt{1-\left(\frac{n_{0}}{2\varpi}\right)^{2}}\,\sqrt{\frac{\pi}{24}}\left(\pm iH_{in}^{\prime}-\frac{\sqrt{e^{2}-1}}{e}H_{in}\right),$
where $n_{0}\equiv m/\Omega$, $\varpi\equiv\pi f/\Omega$, $\gamma\equiv q_{2}/m_{2}-q_{1}/m_{1}$, and $H_{in}^{\prime}=(H_{in-1}-H_{in+1})/2$. The corresponding dipolar scalar memory for $n_{0}=0$ is given by $4\pi R\Delta\phi=\nu M\gamma\hat{\mathbf{k}}\cdot\Delta\mathbf{v}$, with $\Delta\mathbf{v}=-(2a\Omega/e)\mathbf{e}_{X}$ for the hyperbolic orbit. The 0PN dipolar vector waveform can be obtained from the scalar waveform as $\tilde{A}^{0}_{1,\pm 1}=\tilde{\phi}_{1,\pm 1}$, $\tilde{A}^{\text{L}}_{1,\pm 1}=\tilde{\phi}_{1,\pm 1}/v_{g}$ and $\tilde{A}_{1,\pm 1}=\sqrt{2}\,\tilde{\phi}_{1,\pm 1}/v_{g}$, with the replacement $q_{\phi}\to q_{A}$. The spectra are shown in Fig. 27 for different values of $n_{0}$. Even for a massive scalar field, the spectrum in the intermediate-frequency band can mimic a soft waveform for sufficiently small $n_{0}$.
@@
Now we discuss the interferometer response to scalar and vector waves, assuming that the test masses are also described by the worldline action (78) (which are considered, e.g., in [98]). The acceleration of a nonrelativistic test mass induced by weak scalar and vector waves is $\mathbf{F}_{*}\approx\frac{q_{\phi}[\nabla\phi+(\dot{\mathbf{X}}\cdot\nabla\phi+\dot{\phi})\dot{\mathbf{X}}]+q_{A}[-\nabla A^{0}-\dot{\mathbf{A}}+\dot{\mathbf{X}}\times(\nabla\times\mathbf{A})]}{m-q_{\phi}\phi}\approx\frac{q_{\phi}}{m}\nabla\phi-\frac{q_{A}}{m}\left(\nabla A^{0}+\dot{\mathbf{A}}\right)$, where we neglect terms dependent on the test mass velocity $\dot{\mathbf{X}}$. For a plane wave with frequency $\omega=2\pi f$ in the direction $\hat{\mathbf{k}}$, the induced velocity of the $I$-th test mass at $\mathbf{x}_{I}$ is $\mathbf{v}_{I}(t)\equiv\int^{t}dt^{\prime}\,\mathbf{F}_{*}(t^{\prime},\mathbf{x}_{I})\approx-\frac{q_{\phi}}{m}v_{g}\phi\,\hat{\mathbf{k}}+\frac{q_{A}}{m}\left(v_{g}A^{0}\,\hat{\mathbf{k}}-\mathbf{A}\right)$, which leads to the one-way signal:
$y_{rs}(t)\approx-\mathbf{n}_{rs}\cdot[\mathbf{v}_{r}(t)-\mathbf{v}_{s}(t-L)].$ (79)
For a plane wave, the idealized spectrum $\tilde{y}_{rs}(f)$ in the static-detector approximation depends linearly on $\tilde{\phi}$ in the scalar case and on $\{\tilde{A}^{0},\tilde{A}^{\text{L}},\tilde{A}\}$ in the vector case. Soft signals can thus arise from the scalar or vector “velocity memory”, which differ from those due to nonrelativistic waves [98, 97, 95] and, in principle, are also distinguishable from the soft signals of GW memory. The same holds for potentially observable signals from the axion–photon coupling in an axion field $\phi$, for which the one-way signal is independent of $\mathbf{n}_{rs}$ [95].

### 13 Apr 2026 v2

@@ VI Stochastic soft memory signals
A data stream of sufficiently long duration may contain multiple, or even a large number of, events with memory. In the latter case, the detector effectively observes a stochastic GW background (SGWB) of soft memory signals [106]. Even if the individual events cannot be measured, the SGWB may still be detectable if the event rate is sufficiently large. This could be the case, e.g., for the SGWB from compact binary scattering [40, 64]. The SGWB from compact binary mergers also contains a memory contribution, although it is expected to be subleading in the frequency band of interest compared with the oscillatory GW contribution from sources in the same population. In this section, we examine the signal PSD of an idealized SGWB of soft displacement-memory signals at the LISA-like detectors and assess its detectability. We will also briefly comment on the SGWB of soft velocity-memory signals.
@@
$h_{ij}(t,\mathbf{x})=\sum_{\lambda}\int_{-\infty}^{\infty}df\int d\Omega_{2}\,\tilde{h}_{\lambda}(f,\hat{\mathbf{k}})\,e_{ij}^{\lambda}(\hat{\mathbf{k}})\,e^{-2\pi if(t-\hat{\mathbf{k}}\cdot\mathbf{x})}.$
@@
$\left\langle\tilde{h}_{\lambda}(f,\hat{\mathbf{k}})\,\tilde{h}^{*}_{\lambda^{\prime}}(f^{\prime},\hat{\mathbf{k}}^{\prime})\right\rangle=\frac{S_{h}(f)}{16\pi}\,\delta(f-f^{\prime})\,\delta^{2}(\hat{\mathbf{k}},\hat{\mathbf{k}}^{\prime})\,\delta_{\lambda,\lambda^{\prime}},$
@@
For an ensemble of soft displacement-memory signals, the resulting $h_{ij}(t)$ is a superposition of step functions with different step sizes $\mathcal{H}$ and step positions $t_{*}$. Such a time-domain waveform is similar to a Brownian motion and is not wide-sense stationary. Nevertheless, the Welch PSD estimator turns out to be
@@
Now we consider an ensemble of mutually-independent soft displacement-memory signals with a uniform distribution of $\{\tau_{*},\lambda,\sin\beta,\Psi\}$ and a random distribution of $H$. In the present case, we refer to such a SGWB as “idealized”. A single event appears in $h_{33}(t)$ as a step with size $\mathcal{H}=\kappa H$, where $\kappa=-c^{2}_{\beta}\,c_{2\Psi}$. It follows that $\langle\mathcal{H}\rangle=0$, and $\langle\mathcal{H}^{2}\rangle=\sigma_{\kappa}^{2}\left(\langle H\rangle^{2}+\sigma_{H}^{2}\right)$, with $\sigma_{\kappa}^{2}=4/15$. Eqs. (57) and (59) then yield
@@
where $\Delta=\frac{2N}{T}\frac{\langle H\rangle^{2}+\sigma_{H}^{2}}{(2\pi f)^{2}}$. The corresponding energy density $d\rho_{\text{gw}}/df$ [defined by Eq. (55)] is equivalent to a simple sum of the spectral energy flux density $\frac{1}{T}\frac{dE_{\text{gw}}}{R^{2}d\Omega_{2}df}=\frac{H^{2}}{8\pi T}$ associated with each memory jump. For example, if $H$ is distributed uniformly in $0,H_{\text{max}}]$, then $\langle H\rangle^{2}+\sigma_{H}^{2}=\frac{1}{3}H_{\text{max}}^{2}$. To verify this description, a time-domain simulation of the SGWB is performed and the first-generation TDI signal at LISA detector is computed using fastlisaresponse with the realistic ESA orbit. The SGWB is constructed from 4000 sources drawn from a uniform distribution of $\{H,\tau_{*},\lambda,\sin\beta,\Psi\}$. We find that the result agrees well with the analytical prediction to $S_{h}$ and the signal PSD given by Eq. (58), as shown in Fig. 22.
In principle, the signal PSD can be measured [50, 100], and the corresponding squared SNR is given by
@@ VII Summary and discussion
As predicted by general relativity, oscillatory gravitational radiation is generally accompanied by a low-frequency component associated with GW memory. A measurement targeting only the oscillatory part is therefore incomplete. This component, however, can produce soft signals in detectors sensitive to low-frequency bands. At sufficiently low frequencies, such signals take clean, universal forms, with the memory amplitudes as the only intrinsic parameters. We have investigated the prospects for detecting such signals with future space-based interferometers, through observations of individual events and the stochastic background. For a single event, the SNR in LISA, Taiji and TianQin can reach 10 for a displacement memory of amplitude $\sim 10^{-20}$. Such signals could arise from compact-binary scatterings, supernovae, gamma-ray bursts, or certain dark transients (including subsolar-mass BBH mergers). The SNR of a single event in BBO can reach 10 for a displacement memory of amplitude $\sim 10^{-24}$. Such signals are generically expected from stellar-mass compact binary mergers. We examine the parameter-estimation precision for soft displacement-memory signals using simulated Bayesian analyses. The deviation of the GW spectrum from an exact soft waveform can be characterized by a complex correction factor. We examine such corrections in realistic examples of compact binary scattering and quasi-circular BBH mergers, and find that the soft waveform may still serve as an effective template in the presence of relatively large corrections. Many unresolvable soft signals can also form a stochastic gravitational-wave background. For displacement memory, the resulting background is expected to follow a strain power spectral density $S_{h}\propto f^{-2}$. We evaluate the corresponding SNR of the signal power spectrum for the considered detectors. Our results indicate that soft signals from bursts with memory can be promising targets for space-based gravitational-wave detection.
@@ Appendix B Soft signals from dark radiation
As in gravitational radiation, memory can also arise in massless scalar and vector radiation [9]. If the scalar or vector field couples to the test masses in a GW interferometer, this memory can also lead to soft signals (or their mimickers in the case of massive fields), [^12] Note that even in the absence of non-gravitational couplings, dark radiation can still produce GW memory. Perturbatively, the displacement memory is given by [99] $h_{ij}(t,\mathbf{x})\approx\Lambda_{ij,kl}\int dt^{\prime}\,\int d^{3}x^{\prime}\,\frac{4T_{kl}(t^{\prime},\mathbf{x}^{\prime})}{|\mathbf{x}-\mathbf{x}^{\prime}|}\delta(t^{\prime}-t+|\mathbf{x}-\mathbf{x}^{\prime}|)$, where $T_{ij}$ is the energy-momentum tensor of the radiation field. For the massless radiation ($\left\langle T_{ij}(t,\mathbf{x}=R\,\mathbf{n})\right\rangle=\frac{dE}{R^{2}dtd\Omega}n^{i}n^{j}$, where $\frac{dE}{R^{2}dtd\Omega}$ is the angular distribution of the time-averaged energy flux), this leads to Eq. (27) in the distant-source approximation. as we discuss below. Consider a Klein–Gordon field $\phi$ and a Maxwell or Proca field $A_{a}$ sourced by point charges. The worldline action of a point particle with mass $m_{\text{p}}$, scalar charge $q_{\phi}$, and vector charge $q_{A}$ is given by
$S=\int(-m_{\text{p}}+q_{\phi}\phi)\,d\tau+\int q_{A}A_{a}dX^{a},$ (63)
@@
$\phi$ $=\int df\,\sum\tilde{\phi}_{l,m}(f)\,Y_{lm}\,e^{-2\pi ifu},$
$A^{0}$ $=\int df\sum\tilde{A}^{0}_{l,m}(f)\,Y_{lm}\,e^{-2\pi ifu},$
$A^{\text{L}}$ $\equiv\mathbf{A}\cdot\hat{\mathbf{k}}=\int df\sum\tilde{A}_{l,m}^{\text{L}}(f)\,Y_{lm}\,e^{-2\pi ifu},$
$A$ $\equiv\mathbf{A}\cdot(\mathbf{e}_{\Theta}-i\mathbf{e}_{\Phi})=\int df\sum\tilde{A}_{l,m}(f)\,_{-1}Y_{lm}\,e^{-2\pi ifu},$
(see Sec. III for the definition of $\{R,\Theta,\Phi\}$) where $u=t-v_{g}R$, with $v_{g}(f)=\sqrt{1-m_{\text{b}}^{2}/(2\pi f)^{2}}$ being the group velocity ($\text{b}\in\{\phi,A\}$), and $\int_{-\infty}^{\infty}df$ runs only over $2\pi|f|\geq m_{\text{b}}$. The idealized positive-frequency spectra of the non-vanishing modes for the 0PN scalar dipole radiation are
$\frac{\tilde{\phi}_{1,\pm 1}(f)}{\frac{\nu\gamma aM}{R}}=\sqrt{1-\frac{n_{0}^{2}}{n^{2}}}\,\sqrt{\frac{\pi}{24}}\left(\pm iH_{in}^{\prime}-\frac{\sqrt{e^{2}-1}}{e}H_{in}\right),$
where $n_{0}\equiv m_{\text{b}}/\Omega$, $n\equiv 2\pi f/\Omega$, $\gamma\equiv q_{2}/m_{2}-q_{1}/m_{1}$, and $H_{in}^{\prime}=(H_{in-1}-H_{in+1})/2$. Note that $\tilde{\phi}_{1,1}(-f)=-[\tilde{\phi}_{1,-1}(f)]^{*}$. The corresponding dipolar scalar memory for $n_{0}=0$ is given by $4\pi R\Delta\phi=\nu M\gamma\hat{\mathbf{k}}\cdot\Delta\mathbf{v}$, with $\Delta\mathbf{v}=-(2a\Omega/e)\mathbf{e}_{X}$ for the hyperbolic orbit. The 0PN dipolar vector waveform can be obtained from the scalar waveform as $\tilde{A}^{0}_{1,\pm 1}=\tilde{\phi}_{1,\pm 1}$, $\tilde{A}^{\text{L}}_{1,\pm 1}=\tilde{\phi}_{1,\pm 1}/v_{g}$ and $\tilde{A}_{1,\pm 1}=\sqrt{2}\,\tilde{\phi}_{1,\pm 1}/v_{g}$, with the replacement $q_{\phi}\to q_{A}$. The spectra are shown in Fig. 27 for different values of $n_{0}$. Even for a massive scalar field, the spectrum in the intermediate-frequency band can mimic a soft waveform for sufficiently small $n_{0}$.
@@
Now we discuss the interferometer response to scalar and vector waves, assuming that the test masses are also described by the worldline action (63) (which are considered, e.g., in [105]). The acceleration of a nonrelativistic test mass induced by weak scalar and vector waves is $\mathbf{F}_{*}\approx\frac{q_{\phi}}{m_{\text{p}}}\nabla\phi-\frac{q_{A}}{m_{\text{p}}}\left(\nabla A^{0}+\dot{\mathbf{A}}\right)$, where we neglect terms dependent on the test mass velocity $\dot{\mathbf{X}}$. For a monochromatic plane wave with frequency $\omega=2\pi f$ in the direction $\hat{\mathbf{k}}$, the induced velocity of the $I$-th test mass at $\mathbf{x}_{I}$ is $\mathbf{v}_{I}(t)\equiv\int^{t}dt^{\prime}\,\mathbf{F}_{*}(t^{\prime},\mathbf{x}_{I})\approx-\frac{q_{\phi}}{m_{\text{p}}}v_{g}\phi\,\hat{\mathbf{k}}+\frac{q_{A}}{m_{\text{p}}}\left(v_{g}A^{0}\,\hat{\mathbf{k}}-\mathbf{A}\right)$, which leads to the one-way signal:
$y_{rs}(t)\approx-\mathbf{n}_{rs}\cdot[\mathbf{v}_{r}(t)-\mathbf{v}_{s}(t-L_{rs})].$ (64)
For a plane wave, the idealized spectrum $\tilde{y}_{rs}(f)$ in the static-detector approximation depends linearly on $\tilde{\phi}$ in the scalar case and on $\{\tilde{A}^{0},\tilde{A}^{\text{L}},\tilde{A}\}$ in the vector case. Soft signals can thus arise from the scalar or vector “velocity memory”, which differ from those due to nonrelativistic waves [105, 104, 102] and, in principle, are also distinguishable from the soft signals of GW memory. The same holds for potentially observable signals from the axion–photon coupling in an axion field $\phi$, for which the constructible one-way signal is proportional to $\phi(t,\mathbf{x}_{r})-\phi(t-L_{rs},\mathbf{x}_{s})$ [102].